OpenStudy (anonymous):

find the roots of y=2x^2-4x-1

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm 2x^2-4x-1=y}\) \(\large\color{blue}{ \rm 2x^2-4x-1=0}\) \(\large\color{blue}{ \rm 2x^2-4x-1+3=0+3}\) \(\large\color{blue}{ \rm 2x^2-4x+2=3}\) \(\large\color{blue}{ \rm 2(x^2-2x+1)=3}\) \(\large\color{blue}{ \rm x^2-2x+1=3/2}\) \(\large\color{blue}{ \rm (x-1)^2=3/2}\)

OpenStudy (solomonzelman):

can you finish the problem ?

OpenStudy (anonymous):

\[x-1=\sqrt{\frac{ 3 }{ 2 }}\]

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

Yes, but you forgot \(\large\color{blue}{ \rm ± }\) (√2/√3)

OpenStudy (anonymous):

oh, thank you!

OpenStudy (solomonzelman):

this is not it.

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm x-1=± (~\sqrt{3} ~ ~/~\sqrt{2} ~)}\) \(\large\color{blue}{ \rm x=1± (~\sqrt{3} ~ ~/~\sqrt{2} ~)}\) \(\large\color{blue}{ \rm x=1± (~\sqrt{6} ~ ~/2)}\)

OpenStudy (solomonzelman):

there.